# MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM

### v1

@@ Abstract
Large language models (LLMs) have advanced code generation from single-function tasks to competitive-programming problems, but existing multi-agent solutions either rely on costly large-scale (>30B>30B) models or collapse when downsized to small open-source models. We present MapCoder-Lite, which upgrades a single 7B model into four role-specialised agents—retriever, planner, coder, and debugger—using only rank-32, role-specific LoRA adapters ($3\%$ extra parameters). Three lightweight techniques make this possible: (i) trajectory distillation from strong LLMs fixes format fragility in retrieval and debugging, (ii) supervisor-guided correction strengthens planning and coding agents, and (iii) agent-wise LoRA fine-tuning delivers memory-efficient specialisation. Comprehensive evaluation on xCodeEval, APPS, and CodeContests shows that MapCoder-Lite more than doubles xCodeEval accuracy (13.2% → 28.3%), eliminates all format failures, and closes to within six points of a 32B baseline while cutting GPU memory and token-generation time by $4\times$. These results demonstrate that careful agent-wise fine-tuning unleashes high-quality multi-agent coding on a small language model.
MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM
@@ 1 Introduction
LLMs have revolutionized code synthesis, achieving near-perfect accuracy on function-level tasks like HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). Research has now shifted to competitive programming, which demands efficient algorithms, robust implementation, and resilience against hidden test cases. Benchmarks like CodeElo Quan et al. (2025) highlight significant challenges that have spurred the development of massive models, such as OpenAI’s o3 OpenAI et al. (2025) and the 671 B-parameter DeepSeek-V3 DeepSeek-AI et al. (2025b), which dominate leaderboards despite their computational costs and proprietary nature.
To meet the cognitive breadth of competitive programming, recent research has shifted from monolithic prompting to multi-agent code-generation frameworks that allocate distinct roles to specialised language-model agents Islam et al. (2024). These systems separate problem solving into discrete stages—algorithm selection, planning, coding, and debugging—enabling focused expertise and structured information exchange between agents Huang et al. (2024); Hong et al. (2024); Islam et al. (2024, 2025). MapCoder Islam et al. (2024) (Fig. 1) exemplifies this approach with a pipeline where specialized agents handle different aspects of problem solving, narrowing the gap between open-source and proprietary solutions.
However, multi-agent systems present two significant challenges. First, they are computationally intensive when using large models. Our MapCoder implementation with Qwen-2.5-32B on the CodeContests benchmark required 3,000 forward passes, each involving 70 GB of parameter loading, resulting in 27.5 hours of runtime on a single A100 GPU—far exceeding the cost of single-prompt approaches. Second, downsizing the agents to 7B models leads to a sharp decline in reliability (see Fig. 2): (i) format adherence breaks down, with frequent XML schema violations that halt pipeline execution; and (ii) role-specific performance suffers, yielding brittle algorithms, incorrect code, and ineffective debugging. Achieving a balance between competitive performance and computational efficiency remains a key open problem.
@@
  • Trajectory distillation from strong LLMs (Sec. 3.1): For retrieval and debugging roles—where small models often fail—we harvest high-quality traces from Qwen-32B and DeepSeek-V3, filtered through execution tests to ensure format correctness and semantic accuracy.
  • Supervisor-aided cross-agent refinement (Sec. 3.2): For planning and coding roles, we employ a supervisor model that identifies failures, provides targeted feedback, and regenerates problematic steps, creating high-quality training examples without human annotation.
  • Memory-efficient LoRA specialization (Sec. 3.3): All agents share a frozen Qwen-2.5-7B backbone with role-specific low-rank adapters (LoRA Hu et al. (2021)), adding <3% of the parameters of a 7B model while enabling independent fine-tuning.
We conducted a comprehensive evaluation on three representative competitive-programming suites—xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022)—and found that MapCoder-Lite leverages trajectory distillation, supervisor-guided cross-agent refinement, and rank-32 LoRA specialisation to boost xCodeEval accuracy from 13.2% to 28.3%, eliminates every XML-schema failure, and narrows the gap to a 32B MapCoder to within six points, all while cutting GPU memory and token generation time(time per output token) by $4\times$. These results demonstrate that careful agent-wise fine-tuning can unlock high-quality multi-agent code generation on small language models.
2 Background
@@ 2.1 LLMs for Competitive Programming
Large language models have already revolutionised function-level code generation: specialised variants such as Codex, Qwen-2.5-Coder, DeepSeek-Coder, and CodeLlama routinely achieve near-perfect scores on HumanEval, MBPP, and LiveCodeBench Chen et al. (2021); Hui et al. (2024); Guo et al. (2024); Rozière et al. (2024); Jain et al. (2024). But recent community attention has shifted toward the far more demanding arena of competitive programming, where a model must translate a prose problem statement into a complete, time- and memory-efficient program that solves all hidden test cases. Datasets such as APPS Hendrycks et al. (2021), CodeContests Li et al. (2022), and xCodeEval Khan et al. (2024) embody this shift: they assess full-program synthesis, algorithm selection, and constraint reasoning under the stringent standards of contest judging. Progress on these benchmarks now defines the “state of the art” for code-capable LLMs OpenAI et al. (2025); DeepSeek-AI et al. (2025a), but their difficulty also exposes the limits of single-agent prompting, motivating the multi-agent systems discussed next.
@@ 2.2 Multi-Agent Code Generation
As competitive programming problems grow exceptionally in linguistic length and algorithmic subtlety, it is not uncommon for a single LLM prompt to miss hidden edge cases or violate the constraints of the competition. Recent work, therefore, turns to multi-agent pipelines, in which a set of specialised LLM agents collaborate under an explicit workflow Huang et al. (2024); Pan et al. (2025). Among these systems, MapCoder Islam et al. (2024) stands out with its attractive design that performs well in APPS and CodeContests. Islam et al. (2024).
Fig. 1 illustrates MapCoder’s four-stage coding pipeline. A retrieval agent first queries a private corpus of textbook algorithms and returns a short, schema-constrained XML snippet that explains a candidate approach. Next, a planning agent expands this snippet into several step-wise execution plans, and a confidence score is generated for each plan using XML format. Plans are sorted by confidence, and the top one is passed to a coding agent that writes full source code in the required language. Finally, a debugging agent compiles, executes, and patches the program until it clears all unit tests; if it fails, the loop backtracks to the next plan.
Role specialisation makes the pipeline modular: each agent faces a simpler, well-scoped sub-task, so errors can be isolated and corrected step-by-step, yielding higher overall accuracy than a monolithic prompt can achieve. The price of this modularity, however, is that the underlying LLM must possess enough general capability to perform four very different roles—retrieving algorithms, reasoning about plans, writing code, and fixing bugs—without losing coherence across stages.
@@ 2.3 Task-Specific Fine-Tuning
MapCoder relies on large-scale (>30B>30B), general-purpose LLMs at every stage, trading speed and memory for accuracy. Because each agent tackles a well-defined sub-task—retrieving algorithms, planning steps, writing code, or fixing bugs—a natural question is whether a single 7 B backbone can match MapCoder if it is fine-tuned per role. Fine-tuning smaller models has proved effective in other multi-agent settings Zhao et al. (2025); Liang et al. (2025); Shen et al. (2024), and several studies have enhanced code-related abilities through supervised updates or parameter-efficient adapters Fan et al. (2025); Tsai et al. (2024); Yu et al. (2024); Jiang et al. (2024a). To our knowledge, however, no work has comprehensively studied how far role-aligned fine-tuning can push a small LLM inside a multi-agent pipeline for competitive programming.
Model Metric Cost
Number of problems 165
Total runtime 27.53 hours
Qwen32B Input tokens 5.08M tokens
Output tokens 1.60M tokens
Number of API calls 3,095 calls
Table 1: Resource cost of MapCoder using Qwen2.5-32B-Instruct on the CodeContests benchmark.
Figure 2: Representative failure cases in 7B-scale models. (a) Format following error in the retrieval agent: the model fails to generate responses in the required XML structure, leading to format mismatches and processing errors. (b) Incomplete reasoning in the planning agent: a key step is omitted from the solution plan, resulting in failure to fully cover possible cases.
2.4 Challenges
We identify key challenges in multi-agent code generation based on our implementation of MapCoder.
@@ 2.4 Challenges
High Cost with Large Models. We evaluated MapCoder using the Qwen2.5-32B-Instruct model on the CodeContests benchmark, which comprises 165 competitive programming problems. As shown in Table 1, the system required 27.53 hours of runtime, processed approximately 5.08 million input tokens and 1.60 million output tokens, and made 3,095 API calls. This substantial resource usage stems from MapCoder’s multi-agent design involving four distinct agents and multiple iterations for planning and debugging. These results highlight the heavy runtime and memory burden of using large-scale language models (LMs) throughout the pipeline. We hypothesize that replacing each agent with a fine-tuned small LM can significantly reduce token-generation time and GPU memory usage, even with similar token and API call counts.
Failure Cases with Small Models. Fig. 2 illustrates typical failures when downsizing MapCoder’s model to a 7B scale.
  • Format Following Failures: Multi-agent systems often rely on structured outputs in predefined formats. For instance, in MapCoder, the retrieval and planning agents must return XML-formatted responses with tags like <root>, <algorithm> and <confidence>. In Fig. 2(a), the 7B model inserts an extraneous <description> tag and omits the closing </root> tag, violating the expected format. Such structural errors disrupt downstream processing, as subsequent agents or parsers depend on well-formed inputs. These issues are particularly critical in multi-agent pipelines, where format consistency ensures reliable coordination.
  • Low Role Performance: Small models often struggle with role-specific tasks due to limited capacity. As shown in Fig. 2(b), the 32B model’s planning agent enumerates all relevant cases, demonstrating robust and comprehensive reasoning. In contrast, the 7B model omits a key sub-case involving the combination of multiple ticket types. Incomplete planning like this can yield incorrect solutions, even if individual steps seem plausible. Similar shortcomings appear in other roles: the coding agent may generate incorrect logic or indexing, while the debugging agent frequently fails to identify and correct bugs, leading to repeated failures across iterations.
To mitigate these issues, we propose task-specific fine-tuning to improve both format compliance and role execution, thereby enhancing overall code generation accuracy.
@@ 3 Methodology
To replace MapCoder’s 32B backbone with a memory-efficient 7B model while maintaining accuracy, we introduce a role-aligned supervised fine-tuning (SFT) pipeline that tailors one shared 7B backbone to four agents—retrieval, planning, coding, and debugging—via lightweight LoRA adapters. These changes present three specific challenges:
  • Role-specific data scarcity: Public code-generation datasets provide ground-truth programs but lack datasets for retrieval queries, step-wise plans, and debug traces, hindering agent-wise supervision.
  • Cross-agent consistency: Naïvely fine-tuning each agent in isolation overlooks inter-agent dependencies; a fix applied to one step can still fail downstream.
  • Parameter budget inflation: Even if fine-tuning recovers agent-specific capability, storing four independent 7B checkpoints nullifies memory savings, approaching the original 32B footprint.
We address (i) by distilling high-quality trajectories from a strong LLM for retrieval & debugging, tackle (ii) with a supervisor model that diagnoses end-to-end failures and regenerates only the faulty segment for planning & coding, and solve (iii) by sharing a single 7B backbone augmented with rank-32 LoRAs—adding $3\%$ parameters while preserving MapCoder’s modularity.
@@ 3.1 Strong LLM for Retrieval and Debugging
Observations. Two weaknesses repeatedly surface when the 32B backbone is replaced with a 7B model. First, the retrieval agent often violates the mandated XML schema or returns partial algorithm descriptions, so downstream agents cannot even parse its output. Second, the debugging agent fails to patch seemingly simple boundary-case bugs, leaving many programs one step short of success. Experiments with stronger LLMs such as Qwen2.5-32B-Instruct and DeepSeek-V3 show that these larger LLMs seldom make such mistakes, indicating that their trajectories can serve as high-quality supervision for smaller LLMs.
Challenges. Effective role-specific fine-tuning is hampered by two data gaps. (i) Public code-generation benchmarks provide only problem–solution pairs and contain no retrieval queries, step-wise plans, or debug traces; thus, no off-the-shelf corpus targets the intermediate process of the multi-agent coding pipeline. (ii) Even when synthetic trajectories are generated, many are noisy and fail the unit tests, so naïvely training on them would propagate errors rather than correct them.
Methods. Fig. 3 shows the proposed trajectory construction method. Our data pipeline begins by asking strong LLMs—Qwen2.5-32B-Instruct or DeepSeek-V3—to solve each coding task while explicitly printing the intermediate artefacts (i.e. trajectories) obtained from each MapCoder role. In one shot, the model emits an algorithm description for the retrieval agent, a step-wise plan for the planning agent, source code for the coding agent, and a patched version for the debugging agent. Because the same LLM generates the entire trajectory, the four outputs are mutually consistent and naturally aligned with their respective prompts.
To convert these raw trajectories into reliable supervision, we exploit the fact that MapCoder’s final objective—passing all the unit tests—is objectively verifiable. After executing the LLM-produced code, we keep the complete trajectory only if the program passes. This pass‐based filtering eliminates noisy or partially correct traces without any manual inspection: a failure at test time almost always implies that at least one upstream artefact (faulty retrieval cue, flawed plan, or ineffective patch) is wrong, so discarding the whole sequence prevents error propagation during fine-tuning. The surviving set—about 2,200 retrieval examples and 4,200 debugging examples from APPS, xCodeEval, and CodeContests—forms a high-quality, role-aligned corpus that directly trains rank-32 LoRA on the shared 7B backbone. Empirically, models tuned on this filtered dataset regain XML-schema fidelity and bug-repair accuracy comparable to the original 32 B system while retaining the memory footprint advantages of a 7B model (see Table 2 for resulting accuracy enhancement).
@@
Dataset
Pass
Filtering
Dataset
Pass
Filtering
Accuracy (%)
@@
Strong 16.98
Strong Strong 18.87
Strong Supervisor 22.64
Table 2: Ablation study on data source and pass-based filtering for training retrieval, planning, and coding agents on xCodeEval. “Strong” refers to data generated by a large LM (e.g., Qwen2.5-32B, DeepSeek-V3), and “Supervisor” denotes data refined by a supervisor model after failure analysis.

### v2

# MapCoder-Lite: Distilling Multi-Agent Coding into a Single Small LLM

@@ Abstract
Large language models (LLMs) have advanced code generation from single-function tasks to competitive-programming problems, but existing multi-agent solutions either rely on costly large-scale (>30B>30B) models or collapse when downsized to small open-source models. We present MapCoder-Lite, a framework for distilling the complex reasoning of large, multi-agent coding systems into a single 7B model. Our contribution is a novel, three-pillar methodology that synergistically generates, refines, and encodes multi-agent knowledge: (i) pass-based trajectory distillation from strong LLMs fixes format fragility in retrieval and reduces failures in debugging, (ii) supervisor-guided correction with global feedback strengthens planning and coding agents, and (iii) agent-wise LoRA fine-tuning delivers memory-efficient specialisation. Comprehensive evaluation on xCodeEval, APPS, and CodeContests shows that MapCoder-Lite more than doubles xCodeEval accuracy (13.2% → 28.3%), eliminates all format failures, while reducing GPU memory and token-generation time by $4\times$ compared to a 32B model. It also achieves over 10% gains on simpler coding benchmarks, demonstrating broad improvements beyond competitive programming. These results demonstrate that careful agent-wise fine-tuning unleashes high-quality multi-agent coding on a small language model. Our code is publicly available at https://github.com/aiha-lab/MapCoder-Lite.
MapCoder-Lite: Distilling Multi-Agent Coding into a Single Small LLM
@@ 1 Introduction
LLMs have revolutionized code synthesis, achieving near-perfect accuracy on function-level tasks like HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). Research has now shifted to competitive programming, which demands efficient algorithms, robust implementation, and resilience against hidden test cases. Benchmarks like CodeElo Quan et al. (2025) highlight significant challenges that have spurred the development of massive models, such as OpenAI’s o3 OpenAI et al. (2025) and the 671B-parameter DeepSeek-V3 DeepSeek-AI et al. (2025b). These models report strong performance on Codeforces and related benchmarks despite their computational costs and proprietary nature.
Competitive programming involves algorithmically complex problem solving under strict constraints, making it a difficult setting for language models. Single-agent prompting, where one model handles the entire problem-solving process, often falls short Wei et al. (2023); Jiang et al. (2024b); Yasunaga et al. (2024). To overcome this, recent work has explored multi-agent code-generation frameworks that split the task into stages and assign each to a dedicated agent, improving end-to-end performance through role specialization Huang et al. (2024); Hong et al. (2024); Islam et al. (2024, 2025). MapCoder Islam et al. (2024) (Fig. 1) exemplifies this approach, coordinating specialized agents throughout the pipeline.
The effectiveness of multi-agent frameworks typically relies on large-scale (>30B>30B) open models OpenAI et al. (2024); Comanici et al. (2025), which possess the capacity to perform a wide range of specialized roles required across the multi-agent pipeline. Due to their multi-step nature, these frameworks incur significantly higher token usage and API calls than single-agent setups, resulting in increased latency and computational cost.
@@
Their impracticality in resource-constrained settings naturally motivates small-model multi-agent solutions, which align well with the growing industry trend toward on-device AI Gunter et al. (2024); Zhang et al. (2023). In practice, however, deploying small language models (SLMs) under 10B parameters—even those with strong coding abilities Qwen et al. (2025)—often yields limited accuracy gains compared to single-model direct prompting. This performance drop stems from SLMs’ difficulty in following the structured formats (e.g., XML) required for multi-agent communication Xia et al. (2024) and their limited capacity to support the complex reasoning needed across agent roles, leading to failures in retrieval, planning, or debugging.
To make SLM-based multi-agent systems effective, fine-tuning becomes a necessary step. However, this approach poses three key obstacles. First, existing code datasets do not provide intermediate artefacts aligned with the roles in a multi-agent pipeline. Second, fine-tuning each agent independently fails to account for inter-agent dependencies. As illustrated in Fig. 1, later stages rely on the outputs of earlier agents, so errors in one stage can propagate and mislead downstream components. Third, training separate models for each agent increases GPU memory consumption at inference time, undermining the efficiency benefits of using small models in the first place.
@@
  • Pass-based trajectory distillation from strong LLMs (Sect. 5.1): To address the lack of role-specific training data, we collect trajectories from strong LLMs. However, fine-tuning on outputs that are only locally valid often fails to yield correct final solutions. We overcome this by implementing pass-based filtering that exclusively retains trajectories whose final code passes all unit tests.
  • Supervisor-aided cross-agent refinement (Sec. 5.2): We employ a supervisor model that analyzes the full trajectory generated by the small model to detect cross-agent failure patterns and regenerate the responsible agent’s output, guiding the model toward global success and helping bridge the capacity gap.
  • Memory-efficient LoRA specialization (Sec. 5.3): We show that in complex multi-agent settings, LoRA Hu et al. (2021) achieves better accuracy and efficiency than full fine-tuning. This enables all agents to share a frozen Qwen2.5-7B backbone with lightweight, role-specific adapters, adding under 3% extra parameters.
We conducted a comprehensive evaluation on three representative competitive-programming suites—xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022)—and found that MapCoder-Lite leverages trajectory distillation, supervisor-guided cross-agent refinement, and rank-32 LoRA specialisation to boost xCodeEval accuracy from 13.2% to 28.3%, eliminates every XML-schema failure, all while cutting GPU memory and token generation time(time per output token) by $4\times$. These results demonstrate that careful agent-wise fine-tuning can unlock high-quality multi-agent code generation on small language models.
2 Related Work
Figure 2: Representative failure cases of the 7B-scale model across all agents. (a) Retrieval: invalid XML format and incorrect algorithm. (b) Planning: missing key step. (c) Coding: misinterpreted input specification. (d) Debugging: persistent unresolved error. Detailed illustrations are provided in Appendix E.
@@ 2.1 LLMs for Competitive Programming
LLMs have shown strong performance on function-level code generation tasks Hui et al. (2024); Guo et al. (2024); rozière2024codellama, achieving near-perfect scores on HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). Recent work shifts to the harder setting of competitive programming, which requires generating full, efficient programs that pass hidden tests. Benchmarks like APPS Hendrycks et al. (2021), CodeContests Li et al. (2022), and xCodeEval Khan et al. (2024) capture this challenge and now define the state of the art OpenAI et al. (2025); DeepSeek-AI et al. (2025a), motivating the multi-agent systems that follow.
@@ 2.2 Multi-Agent Code Generation
Recent work has proposed multi-agent pipelines to better handle the linguistic complexity and algorithmic subtlety of competitive programming, which often cause single prompts to miss edge cases or violate constraints Huang et al. (2024); Pan et al. (2025). Among these, MapCoderIslam et al. (2024) adopts a four-stage pipeline (Fig.1) that performs well on APPS and CodeContests. A retrieval agent first selects an algorithm from a private corpus and returns a schema-constrained XML snippet. A planning agent then expands this into step-wise plans with confidence scores. The top plan is passed to a coding agent, which generates the full source code. Finally, a debugging agent tests and patches the code until it passes all unit tests, backtracking to alternative plans if needed. This modular design improves accuracy via role specialization, but also demands broad capability from the underlying LLM across all stages.
@@ 2.3 Task-Specific Fine-Tuning
MapCoder uses strong LLMs at each stage, trading efficiency for accuracy. A natural question is whether a single small model (<10B), fine-tuned per role, can achieve similar results. While prior work has shown the value of fine-tuning in multi-agent Zhao et al. (2025); Liang et al. (2025); Shen et al. (2024) and code tasks Fan et al. (2025); Tsai et al. (2024); Yu et al. (2024); Jiang et al. (2024a), no study has systematically explored role-aligned fine-tuning for small LLMs in competitive programming pipelines. To our knowledge, however, no work has comprehensively studied how far role-aligned fine-tuning can push a small LLM inside a multi-agent pipeline for competitive programming.
3 Analysis of Multi-Agent Limitations
3.1 High Cost with Large Models
@@
We evaluated MapCoder using the Qwen2.5-32B-Instruct model on the CodeContests benchmark, which comprises 165 competitive programming problems. The system required 27.53 hours of runtime, processed approximately 5.08 million input tokens and 1.60 million output tokens, and made 3,095 API calls. This substantial resource usage stems from MapCoder’s multi-agent design involving four distinct agents and multiple iterations for planning and debugging. These results highlight the heavy runtime and memory burden of using large-scale language models throughout the pipeline. We therefore hypothesize that replacing each agent with a fine-tuned SLM can substantially reduce token-generation time and GPU memory usage, even under comparable token and API call counts.
3.2 Failure Cases of Small Models
Format Following Failures. Multi-agent systems often rely on structured outputs in predefined formats Yang et al. (2025); Tang et al. (2025). In MapCoder, for example, retrieval and planning agents are required to produce XML-formatted responses such as <root>, <algorithm>, and <confidence> for downstream parsing. However, small models frequently violate these schemas—omitting or misplacing tags—which disrupts subsequent processing and halts pipeline execution (Fig. 2a). This issue is amplified by the weaker format-following ability of open-source SLMs compared to proprietary LLMs Xia et al. (2024), underscoring the importance of strict structural adherence in multi-agent workflows.
Low Role Performance. Small models often struggle with role-specific tasks due to limited capacity. Fig. 2 illustrates representative failures across agents: (a) the retrieval agent produces incorrect or misleading algorithm descriptions, corrupting shared context for all downstream stages; (b) the planning agent outputs superficially correct but incomplete plans, often missing subtle edge cases; (c) the coding agent introduces logical or I/O errors even when given valid plans, yielding incorrect or unexecutable code; and (d) the debugging agent fails to detect or fix simple bugs, resulting in repeated ineffective patches. Together, these failures indicate that, without stronger supervision or greater capacity, individual agents act as bottlenecks that undermine the reliability of the entire pipeline.
4 Challenges
The multi-agent approach provides a parameter-free mechanism for orchestrating specialized roles via prompting, yet SLMs consistently underperform in this zero-shot setting. This suggests that prompt-based assignment alone is insufficient for SLMs, necessitating a shift toward parameter-driven optimization Du et al. (2025). However, effectively fine-tuning SLMs within a multi-agent framework presents several non-trivial challenges.
4.1 Lack of Role-Specific Training Data.
Multi-agent fine-tuning requires high-quality intermediate supervision tailored to each agent’s role. Existing approaches rely either on reconstructing intermediate signals from public datasets Shen et al. (2024) or on self-collection using SLM-generated outputs Liang et al. (2025). However, public code benchmarks are not designed for multi-agent pipelines and lack role-aligned artefacts, while self-collected data from small models is often malformed or incomplete due to limited capacity and task complexity. This underscores the need to generate new, role-specific supervision tailored to multi-agent training.
4.2 Limited Global Awareness.
In multi-agent workflows, information flows sequentially, making end-to-end success reliant on the coherence of intermediate outputs across stages. Yet agents trained in isolation lack awareness of these dependencies, leading to local errors that propagate and compromise the final outcome. For instance, when the final code fails, it is often unclear whether the root cause lies in a flawed plan from the planning agent or in an incorrect implementation by the coding agent, even if each agent appears to perform its role plausibly. Such failures show that success depends not only on individual competence but also on aligned interactions. Without global awareness, even well-tuned agents may fall short.
4.3 Inefficiency of Full Model Fine-Tuning
Previous multi-agent fine-tuning approaches typically rely on OpenAI’s fine-tuning API Zhao et al. (2025) or full-parameter tuning for each agent Shen et al. (2024); Zeng et al. (2023). However, full fine-tuning scales memory usage linearly with the number of agents, negating one of the primary advantages of SLMs—their efficiency in memory and deployment. Furthermore, full fine-tuning does not necessarily guarantee superior performance over parameter-efficient methods such as LoRA Hu et al. (2021).
@@ 5 Methodology
To overcome the limitations of applying multi-agent code generation to SLMs, we propose a role-aligned supervised fine-tuning pipeline that equips a single 7B backbone with specialized behavior for four agents via lightweight LoRA adapters. This design directly addresses the three main challenges identified in our analysis:
  • Lack of role-specific data. To compensate for the absence of intermediate supervision, we distill high-quality trajectories from strong LLMs, using execution results to retain only clean, agent-specific samples (Sec. 5.1).
  • Limited global awareness. To maintain cross-agent consistency, we introduce a supervisor-guided refinement mechanism that detects failures and regenerates only the faulty component, yielding coherent, context-aware training data (Sec. 5.2).
  • Inefficiency of full-model fine-tuning. We address parameter and memory overhead by applying rank-32 LoRA adapters to a shared frozen 7B backbone, enabling agent-wise specialization with minimal additional cost (Sec. 5.3).
Together, these techniques allow us to retain the advantages of the multi-agent approach while making it viable for deployment on open-source, resource-efficient language models.
@@ 5.1 Strong LLM for Retrieval and Debugging
Fig. 3 shows the proposed trajectory construction method. Our data pipeline begins by asking strong LLMs, namely Qwen2.5-32B-Instruct and DeepSeek-V3, to solve each coding task while explicitly printing intermediate artefacts (i.e., trajectories) produced by each MapCoder role.
To build reliable training data, we collect trajectories from strong LLMs and keep only those whose final code passes all unit tests. Unlike rejection sampling based solely on local validity at the single-agent level Zelikman et al. (2022); Zeng et al. (2023), our pass-based filtering ensures end-to-end success across all roles. This focuses fine-tuning on trajectories verified through full execution, yielding accuracy gains over locally valid samples (Table 1, row 2 vs. row 3).
@@
Dataset Filtering Dataset Filtering Accuracy (%)
@@
Strong Format 16.98
Strong Pass Strong Pass 18.87
Strong Pass Supervisor Pass 22.64
Table 1: Ablation study on data source and filtering methods for training retrieval, planning, and coding agents on xCodeEval. “Format” filtering keeps samples with valid single-agent outputs, while “Pass” retains only those where the final program passes all unit tests. “Strong” denotes data from a large LLM; “Supervisor” indicates trajectories refined after failure analysis.
Figure 4: Cosine similarity between algorithm descriptions generated by base and fine-tuned models and ground-truth algorithm tags for 50 xCodeEval problems. Darker cells indicate stronger alignment.
Our analysis further shows that self-collected trajectories from a 7B model often mislabel tasks, such as overpredicting dynamic programming, and exhibit limited algorithmic diversity. As a result, fine-tuning on these traces leads to lower performance (7BFT). In contrast, trajectories generated by 32B align more closely with ground-truth tags and provide more accurate algorithm descriptions. Fine-tuning on these trajectories (7BFT(32B)) significantly improves performance, demonstrating that strong-LLM supervision transfers both correctness and algorithmic diversity to smaller models.